Solve 6x+x(x-3)=0
Find the x-intercepts of f(x) = 6x+x(x-3)
Solutions in ordered pair.
x-intercept in ordered pair.

Answer by edjones(8007)

6x+x(x-3)=0
6x+x^2-3x=0
x^2+3x=0
x(x+3)=0
x=0, x=-3
The x intercepts are also 0 and -3.
(0, -3) ordered pair.
.
Ed